# A Systematic Literature Review on Vehicular Collaborative Perception – A Computer Vision Perspective

### v1

@@ VI-D Collaborative Motion Prediction
Collaborative motion prediction leverages the collective intelligence of multiple observing agents, integrating diverse data sources to enhance the accuracy and robustness of predictions. This cooperative approach not only mitigates the effects of individual sensor occlusions but also provides a more reliable prediction framework compared to isolated mechanisms. Motion prediction can be descirbed as forecasting the trajectory of bounding boxes or forcasting the BEV map. All papers on Collaborative Motion Prediction (CMP) that meet the selection criteria are summarized in Table X.
TABLE X: Overview of the method for collaborative motion prediction (CMP). V: Vehicle, I: Infrastructure, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion, Graph: Graph-based Fusion.
@@ VI-E Collaborative Lane Detection
Lane detection is a critical component for Advanced Driver Assistance Systems (ADAS) and automated driving (AD), as it provides essential information for path planning and vehicle control. High-definition Map (HD Map), though effective, is expensive to create, maintain, and scale. This makes real-time lane detection and online HD Map learning increasingly important. However, lane detection, like other perception tasks, faces challenges such as visual occlusion and limited perception range, particularly in urban intersections with dense traffic, where multi-agent collaboration offers a potential solution. This section categorizes collaborative lane detection into two main approaches: curve-model-based methods and BEV-map-based methods. Lane information can be represented using curve models, which are more data-efficient and require less bandwidth, or BEV segmentation, which provides pixel-level detail with higher resolution and greater robustness to noise. While both approaches offer substantial potential, they remain under-explored and require deeper investigation. All papers on Collaborative Lane Detection (CLD) that meet the selection criteria are summarized in Table XI.
TABLE XI: Overview of the methods for CLD. V: Vehicle, Trad Feat: Traditional Feature Fusion.
@@ VI-F Multi-Task and Task-Agnostic
Autonomous vehicle navigation requires addressing various perception tasks, from object detection to semantic segmentation. Traditionally, these tasks are performed independently, consuming significant computational resources. To optimize resource usage and enhance performance across multiple tasks simultaneously, researchers have proposed multi-task learning pipelines to address multiple perception tasks. All papers on multi-task and task-agnostic method that meet the selection criteria are summarized in Table XII. For example, V2XFormer [103] simultaneously performs object detection, motion prediction, and accident prediction. Similarly, CoBEVT [94] handles both object detection and semantic segmentation in parallel. V2VNet [84] is also able to conduct object detection and motion prediction at the same time.
TABLE XII: Overview of methods for multi-task pipeline and task-agnostic pipeline. OD: Object detection, OT: Object tracking, MP: Motion prediction, AP: Accident prediction, SS: Semantic segmentation, V: Vehicle, I: Infrastructure, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion, Graph: Graph-based Fusion..
@@
BEV-V2X[102] IEEE T-IV 2023 Agnostic V, I Intermediate Atten Feat SS,MP
@@
FF-Tracking[99] CVPR 2023 LiDAR, Camera V, I Intermediate Trad Feat OD,OT
CoBEVT[94] CoRL 2023 Camera V Intermediate Trad Feat OD,SS
@@ VII Approaches to address realistic issues (RQ1)
In the initial stages of research, the focus on CP primarily focused on the collaboration process and fusion strategies under ideal conditions, often relying on unrealistic assumptions such as precise localization and ideal communication conditions. However, CP algorithms encounter numerous challenges when applied in real-world scenarios. This section summarizes these practical issues and their corresponding solutions.
@@ VII-A Localization Errors
Accurate spatial alignment is essential for effective data fusion among different agents. However, errors in localization can lead to data misalignment, significantly impacting perception accuracy. To tackle this issue, researchers focus on correcting the relative pose before alignment [8, 81, 18, 102, 51, 90, 87, 15, 55, 61], the process as shown in Figure 12. Approaches to address localization errors are summarized in Table XIII. The various approaches to address this problem can be categorized into three levels: raw-sensor, object, and feature levels. These methods are comparatively analyzed below.
TABLE XIII: Overview of the methods for addressing pose error.
@@ VII-B Time Latency
Effective data fusion in CAVs necessitates that data be temporally aligned. However, achieving this alignment is challenging due to several factors. Communication delays, interruptions, varying processing times, and differing data rates across vehicles can all lead to temporal misalignment. To mitigate these issues, three levels of data are utilized: object-level data, feature-level data, and occupancy-level data.
@@ VII-B1 Object-Level
TABLE XIV: Overview of methods for addressing latency at the feature level. V: Vehicle, I: Infrastructure.
@@
FF-Tracking[99] CVPR 2023 LiDAR, Camera V,I Feature flow prediction
@@ VII-B2 Feature-Level
Feature-level approaches have gained significant attention in addressing temporal alignment challenges for intermediate collaboration in CAVs. The overview of feature-level approaches to address time latency is shown in Table XIV. One notable approach is the latency-aware collaborative perception system introduced by Lei et al. [19]. This system employs SyncNet, a latency compensation module utilizing Long Short-Term Memory (LSTM) networks to estimate real-time features for collaboration. SyncNet has demonstrated effectiveness in enhancing intermediate collaboration, particularly in high-latency scenarios. However, real-time feature prediction can be computationally intensive. An alternative method, proposed by Wei et al. [48], focuses on BEV flow prediction within the CP framework. This approach, called CoBEVFlow, generates spatial regions of interest (ROIs) based on received perceptual feature maps. By associating correlated ROIs across message sequences, it calculates motion vectors and estimates object positions at specific timestamps. The resulting BEV flow map is used to adjust the spatial position of features, ensuring temporal alignment with ego features for efficient aggregation. Yu et al. [115] introduce another technique called Feature Flow Net (FFNet), which employs feature flow prediction. This method describes feature changes over time, enabling direct prediction of aligned features at the current timestamp of collaborating vehicles. Similarly, the How2comm [52] framework utilizes feature flow prediction but refines it with a scale matrix. This scaling of predicted features has been shown to enhance temporal alignment effectiveness.
@@ VII-C Communication Bandwidth Constraints
In any system requiring communication, bandwidth can become a bottleneck when multiple entities participate and actively contribute to sharing data. In CP, several entities (vehicles or infrastructure) collect, share, and aggregate perception data. In Europe, the European Telecommunications Standards Institute (ETSI) has standardized assumptions over the vehicular network that need to be considered when designing CP algorithms. The vehicular network is ad hoc, participants establish communication in a self-organizing manner, and safety applications, such as CP, generate messages within the limits of a Packet Data Unit (PDU). The PDU size is defined by the access layer protocol, meaning there are bandwidth and single transmission constraints. This section discusses approaches that alleviate the data volume required by CP algorithms through data selection (filtering), data compression, and cooperator selection. Table XV presents an overview of methods to address communication bandwidth constraints.
TABLE XV: Overview of methods for addressing communication efficiency. V: Vehicle, I: Infrastructure.
@@ VII-E Domain Shifts
CP frameworks for CAVs face significant challenges due to domain shift, a problem often under-explored in the field. This section examines the approaches to address different types of domain shift caused by training data, sensor characteristics, and the transition from simulation to real-world environments (Sim2Real). An overview can be seen in Table XVI.
TABLE XVI: Overview of methods for addressing domain shift. V: Vehicle, I: Infrastructure.
@@ VII-F Heterogeneity
Heterogeneity within CP systems presents a significant challenge, primarily caused by differences in sensors and perception models across agents. CAVs on the road are often manufactured by various companies, leading to differences in sensor types and data processing models across vehicles from different Original Equipment Manufacturers (OEMs). This section provides a summary of the approaches used to address both model heterogeneity and modality heterogeneity within CP systems, as listed in Table XVII.
TABLE XVII: Overview of methods for addressing the problem of heterogeneity. V: Vehicle, I: Infrastructure.
@@ VIII Evaluation Methods for Collaborative Perception (RQ2-4)
Evaluation methods are a critical aspect of research on CP, complementing the development of CP approaches. This section provides an overview of the current evaluation methodologies employed in the surveyed studies. Figure 13 illustrates the distribution of surveyed studies across different evaluation methods. Synthetic datasets are the most commonly used, with 67 studies relying on this approach. In contrast, only 3 studies perform evaluations through real-world experiments. Section VIII-A describes the evaluation methodologies in detail, while Section VIII-B focuses on the evaluation scenarios. Additionally, Section VIII-C presents the metrics and ablation studies used to assess CP approaches.
Figure 13: Number of surveyed studies per evaluation method. Real. Data.: Real-world dataset, Syn. Data.: Synthetic dataset, Real. Exp.: Real-world experiments, Sim. Exp.: Simulation experiments.
@@ VIII-A1 Real World Datasets
Table XVIII provides a summary of publicly available datasets for CP. These datasets predominantly focus on vehicle-to-infrastructure (V2I) collaboration, with limited attention given to vehicle-to-vehicle (V2V) interactions. The absence of datasets that integrate both V2I and V2V collaborations highlights a significant gap in existing resources, underscoring the need for more comprehensive datasets to advance CP research.
@@
Table XIX delves deeper into V2X configurations, examining critical aspects such as the number of connected vehicles, localization methods, and time synchronization protocols. Most datasets involve three or fewer CAVs, as exemplified by LUCOOP [123]. Time synchronization is generally asynchronous, with a maximum latency of 50 milliseconds between entities. To ensure accurate ground-truth localization, hybrid localization approaches are commonly employed, combining multiple techniques such as HD Map and Real Time Kinematic (RTK) to minimize positional errors. These precise localization methods play a pivotal role in enhancing CP system performance and fostering effective collaboration between vehicles and infrastructure.
TABLE XVIII: Overview of all publicly available real world datasets for CP that include infrastructure perspective. L indicates LiDAR and RGB denotes camera sensor in the modalities. For the tasks the datasets include object detection (OD) – 3D if not indicated otherwise –, object tracking (OT), motion prediction (MP) and domain adaption (DA).
@@
TABLE XIX: Additional technical information for the publicly available real world datasets. A “-” indicates the absence of information in the referenced publication.
@@
DAIR-V2X-C [121] 1 Hybrid async (10$\sim$30ms) + sync (<10ms)
@@
DAIR-V2X-Seq [99] 1 Hybrid async + sync (<10ms)
@@ VIII-A2 Synthetic Datasets
Table XX summarizes the synthetic datasets used in CP. These datasets are predominantly generated using CARLA, often paired with frameworks like OpenCDA [126], which integrates CARLA with SUMO for traffic simulation. The use of simulation significantly reduces the effort required to create scenarios involving multiple CAVs compared to real-world settings. Unlike real-world datasets, synthetic datasets frequently support both V2V and V2I communication, making them highly versatile for CP research. Additionally, many synthetic datasets enable cooperation among more than three CAVs, further enhancing their applicability. These datasets also exhibit greater diversity in sensor modalities, incorporating LiDAR and RGB cameras, with some extending to include depth information.
@@
A significant characteristic of synthetic datasets is their reliance on idealized system conditions. As outlined in Table XXI, most datasets assume perfect time synchronization between connected entities, with the exception of [48]. Additionally, simulators provide precise ground-truth localization, resulting in error-free localization performance. While these conditions simplify evaluation, they may not fully reflect the challenges of real-world scenarios.
@@
TABLE XX: General overview of synthetic datasets for CP. L indicates LiDAR as modality. The datasets support object detection (OD), object tracking (OT), semantic segmentation (SS), motion prediction (MP) and accident prediction (AP).
@@
CODD [44] 2021 V2V L OD 2 13k CARLA Scenario
V2X-Sim1.0 [45] 2021 V2V L & RGBD OD 3 10k CARLA & Sumo Scenario
OPV2V [127] 2022 V2V L & RGB OD 1 12k OpenCDA Scenario
V2X-Sim 2.0 [128] 2022 V2V & V2I L & RGBD OD, OT, SS 3 10k CARLA & Sumo Scenario
V2XSet [16] 2022 V2V & V2I L OD 1 12k OpenCDA Scenario
IRV2V [48] 2023 V2V L & RGB OD NA 8k CARLA Scenario
OPV2V-H [77] 2024 V2V L & RGBD OD 1 10k OpenCDA Scenario
Semantic OPV2V [96] 2024 V2V L & RGBD SS 12 10k OpenCDA Scenario
DeepAccident [103] 2024 V2V & V2I L & RGB OD, OT, SS, MP, AP 2 57k CARLA Weather, daytime
TABLE XXI: Overview of more technical aspects of the synthetic datasets. Default synchronization means perfect synchronization

### v3

@@ VI-D Collaborative Motion Prediction
Collaborative motion prediction leverages the collective intelligence of multiple observing agents, integrating diverse data sources to enhance the accuracy and robustness of predictions. This cooperative approach not only mitigates the effects of individual sensor occlusions but also provides a more reliable prediction framework compared to isolated mechanisms. Motion prediction can be descirbed as forecasting the trajectory of bounding boxes or forcasting the BEV map. All papers on Collaborative Motion Prediction (CMP) that meet the selection criteria are summarized in Table XI.
TABLE XI: Overview of the method for collaborative motion prediction (CMP). V: Vehicle, I: Infrastructure, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion, Graph: Graph-based Fusion.
@@ VI-E Collaborative Lane Detection
Lane detection is a critical component for Advanced Driver Assistance Systems (ADAS) and automated driving (AD), as it provides essential information for path planning and vehicle control. High-definition Map (HD Map), though effective, is expensive to create, maintain, and scale. This makes real-time lane detection and online HD Map learning increasingly important. However, lane detection, like other perception tasks, faces challenges such as visual occlusion and limited perception range, particularly in urban intersections with dense traffic, where multi-agent collaboration offers a potential solution. This section categorizes collaborative lane detection into two main approaches: curve-model-based methods and BEV-map-based methods. Lane information can be represented using curve models, which are more data-efficient and require less bandwidth, or BEV segmentation, which provides pixel-level detail with higher resolution and greater robustness to noise. While both approaches offer substantial potential, they remain under-explored and require deeper investigation. All papers on Collaborative Lane Detection (CLD) that meet the selection criteria are summarized in Table XII.
TABLE XII: Overview of the methods for CLD. V: Vehicle, Trad Feat: Traditional Feature Fusion.
@@ VI-F Multi-Task and Task-Agnostic
Autonomous vehicle navigation requires addressing various perception tasks, from object detection to semantic segmentation. Traditionally, these tasks are performed independently, consuming significant computational resources. To optimize resource usage and enhance performance across multiple tasks simultaneously, researchers have proposed multi-task learning pipelines to address multiple perception tasks. All papers on multi-task and task-agnostic method that meet the selection criteria are summarized in Table XIII. For example, V2XFormer [103] simultaneously performs object detection, motion prediction, and accident prediction. Similarly, CoBEVT [94] handles both object detection and semantic segmentation in parallel. V2VNet [84] is also able to conduct object detection and motion prediction at the same time.
TABLE XIII: Overview of methods for multi-task pipeline and task-agnostic pipeline. OD: Object detection, OT: Object tracking, MP: Motion prediction, AP: Accident prediction, $S$: Semantic segmentation, V: Vehicle, I: Infrastructure, Raw: Raw data fusion, Trad Feat: Traditional Feature Fusion, Atten Feat: Attention Feature Fusion, Obj Fusion: Object-level Fusion, Graph: Graph-based Fusion..
@@
BEV-V2X[102] IEEE T-IV 2023 Agnostic V, I Intermediate Atten Feat $S$,MP
@@
$F$-Tracking[99] CVPR 2023 LiDAR, Camera V, I Intermediate Trad Feat OD,OT
CoBEVT[94] CoRL 2023 Camera V Intermediate Trad Feat OD,$S$
@@ VII Approaches to address realistic issues (RQ1)
In the initial stages of research, the focus on CP primarily focused on the collaboration process and fusion strategies under ideal conditions, often relying on unrealistic assumptions such as precise localization and ideal communication conditions. However, CP algorithms encounter numerous challenges when deployed in real-world scenarios. This section summarizes these practical issues and their corresponding solutions.
@@ VII-A Localization Errors
Accurate spatial alignment is essential for effective data fusion among different agents. However, errors in localization can lead to data misalignment, significantly impacting perception accuracy. To tackle this issue, researchers focus on correcting the relative pose before alignment [8, 81, 18, 102, 51, 90, 87, 15, 55, 61], the process as shown in Figure 8. Approaches to address localization errors are summarized in Table XIV. The various approaches to address this problem can be categorized into three levels: raw-sensor, object, and feature levels. These methods are comparatively analyzed below.
TABLE XIV: Overview of the methods for addressing pose error.
@@ VII-B Time Latency
Effective data fusion in CAVs necessitates that data be temporally aligned. In practice, Connected Autonomous Vehicle (CAV) s synchronize their clocks with Coordinated Universal Time (UTC) primarily through Global Navigation Satellite System (GNSS) signals, while the Network Time Protocol (NTP) may serve as a fallback when GNSS is unavailable. Achieving perfect alignment, however, remains challenging due to factors such as communication delays, interruptions, heterogeneous processing times, and varying data rates across vehicles, all of which can introduce temporal misalignment. To mitigate these issues, three levels of data are utilized: object-level data, feature-level data, and occupancy-level data.
@@ VII-B1 Object-Level
TABLE XV: Overview of methods for addressing latency at the feature level. V: Vehicle, I: Infrastructure.
@@
$F$-Tracking[99] CVPR 2023 LiDAR, Camera V,I Feature flow prediction
@@ VII-B2 Feature-Level
Feature-level approaches have gained significant attention in addressing temporal alignment challenges for intermediate collaboration in CAVs. The overview of feature-level approaches to address time latency is shown in Table XV. One notable approach is the latency-aware collaborative perception system introduced by Lei et al. [19]. This system employs SyncNet, a latency compensation module utilizing Long Short-Term Memory (LSTM) networks to estimate real-time features for collaboration. SyncNet has demonstrated effectiveness in enhancing intermediate collaboration, particularly in high-latency scenarios. However, real-time feature prediction can be computationally intensive. An alternative method, proposed by Wei et al. [48], focuses on BEV flow prediction within the CP framework. This approach, called CoBEVFlow, generates spatial regions of interest (ROIs) based on received perceptual feature maps. By associating correlated ROIs across message sequences, it calculates motion vectors and estimates object positions at specific timestamps. The resulting BEV flow map is used to adjust the spatial position of features, ensuring temporal alignment with ego features for efficient aggregation. Yu et al. [115] introduce another technique called Feature Flow Net (FFNet), which employs feature flow prediction. This method describes feature changes over time, enabling direct prediction of aligned features at the current timestamp of collaborating vehicles. Similarly, the How2comm [52] framework utilizes feature flow prediction but refines it with a scale matrix. This scaling of predicted features has been shown to enhance temporal alignment effectiveness.
@@ VII-C Communication Bandwidth Constraints
In any system requiring communication, bandwidth can become a bottleneck when multiple entities participate and actively contribute to sharing data. In CP, several entities (vehicles or infrastructure) collect, share, and aggregate perception data. In Europe, the European Telecommunications Standards Institute (ETSI) has specified functional requirements for collective awareness and cooperative perception applications. These specifications establish well-recognized networking constraints that must be considered in the design of CP algorithms. The vehicular network is ad hoc, participants establish communication in a self-organizing manner, and safety applications, such as CP, generate messages within the limits of a Packet Data Unit (PDU). The PDU size is defined by the access layer protocol, which imposes bandwidth limitations and the requirement that safety messages such as Cooperative Awareness Message (CAM) and Collective Perception Message (CPM) must fit into a single PDU, as they are broadcast only once without retransmission or forwarding.
In practice, the bandwidth available for vehicular communication is highly constrained. For instance, IEEE 802.11p/DSRC provides a theoretical data rate of up to 27 Mbps in a 10 MHz channel, but the effective throughput in dense traffic environments is typically below 10 Mbps due to protocol overhead and channel contention [116, 117]. Similarly, LTE-V2X operating in a 10 MHz channel can achieve around 15 Mbps under ideal conditions, yet its capacity decreases substantially as the number of vehicles increases [118]. These limitations render the direct transmission of raw sensor data, such as LiDAR point clouds or high-resolution camera frames, largely infeasible in real-world deployments. This motivates the need for more efficient communication strategies that focus on transmitting intermediate features, or final perception outputs rather than raw data.
To mitigate these constraints, recent research has explored approaches such as data selection to filter transmitted information, data compression to reduce message size, and cooperator selection to restrict the number of entities participating in the CP algorithm. Table XVI provides an overview of representative methods that address communication bandwidth challenges in CP.
TABLE XVI: Overview of methods for addressing communication efficiency. V: Vehicle, I: Infrastructure.
@@ VII-E Domain Shifts
CP frameworks for CAVs face significant challenges due to domain shift, a problem often under-explored in the field. This section examines the approaches to address different types of domain shift caused by training data, sensor characteristics, and the transition from simulation to real-world environments (Sim2Real). An overview can be seen in Table XVII.
TABLE XVII: Overview of methods for addressing domain shift. V: Vehicle, I: Infrastructure.
@@ VII-F Heterogeneity
Heterogeneity within CP systems presents a significant challenge, primarily caused by differences in sensors and perception models across agents. CAVs on the road are often manufactured by various companies, leading to differences in sensor types and data processing models across vehicles from different Original Equipment Manufacturers (OEMs). This section provides a summary of the approaches used to address both model heterogeneity and modality heterogeneity within CP systems, as listed in Table XVIII.
TABLE XVIII: Overview of methods for addressing the problem of heterogeneity. V: Vehicle, I: Infrastructure.
@@ VIII Evaluation Methods for Collaborative Perception (RQ2-4)
Evaluation methods are a critical aspect of research on CP, complementing the development of CP approaches. This section provides an overview of the current evaluation methodologies employed in the surveyed studies. Section VIII-A describes the evaluation methodologies in detail, while Section VIII-B focuses on the evaluation scenarios. Additionally, Section VIII-C presents the metrics and ablation studies used to assess CP approaches.
@@ VIII-A1 Real World Datasets
Table XIX provides a summary of publicly available datasets for CP. These datasets predominantly focus on vehicle-to-infrastructure (V2I) collaboration, with limited attention given to vehicle-to-vehicle (V2V) interactions. The absence of datasets that integrate both V2I and V2V collaborations highlights a significant gap in existing resources, underscoring the need for more comprehensive datasets to advance CP research.
@@
Table XX delves deeper into V2X configurations, examining critical aspects such as the number of connected vehicles, localization methods, and time synchronization protocols. Most datasets involve three or fewer CAVs, as exemplified by LUCOOP [126]. Time synchronization is generally asynchronous, with a maximum latency of 50 milliseconds between entities. To ensure accurate ground-truth localization, hybrid localization approaches are commonly employed, combining multiple techniques such as HD Map and Real Time Kinematic (RTK) to minimize positional errors. These precise localization methods play a pivotal role in enhancing CP system performance and fostering effective collaboration between vehicles and infrastructure.
TABLE XIX: Overview of all publicly available real world datasets for CP that include infrastructure perspective. L indicates LiDAR and RGB denotes camera sensor in the modalities. For the tasks the datasets include object detection (OD) – 3D if not indicated otherwise –, object tracking (OT), motion prediction (MP) and domain adaption (DA).
@@
TABLE $X$: Additional technical information for the publicly available real world datasets. A “-” indicates the absence of information in the referenced publication.
@@
DAIR-V2X-C [124] 1 Hybrid
[…]
+ sync (<10ms)
@@
DAIR-V2X-Seq [99] 1 Hybrid
async +
sync (<10ms)
@@ VIII-A2 Synthetic Datasets
Table XXI summarizes the synthetic datasets used in CP. These datasets are predominantly generated using CARLA, often paired with frameworks like OpenCDA [129], which integrates CARLA with SUMO for traffic simulation. The use of simulation significantly reduces the effort required to create scenarios involving multiple CAVs compared to real-world settings. Unlike real-world datasets, synthetic datasets frequently support both V2V and V2I communication, making them highly versatile for CP research. Additionally, many synthetic datasets enable cooperation among more than three CAVs, further enhancing their applicability. These datasets also exhibit greater diversity in sensor modalities, incorporating LiDAR and RGB cameras, with some extending to include depth information.
@@
A significant characteristic of synthetic datasets is their reliance on idealized system conditions. As outlined in Table XXII, most datasets assume perfect time synchronization between connected entities, with the exception of [48]. Additionally, simulators provide precise ground-truth localization, resulting in error-free localization performance. While these conditions simplify evaluation, they may not fully reflect the challenges of real-world scenarios.
@@
TABLE XXI: General overview of synthetic datasets for CP. L indicates LiDAR as modality. The datasets support object detection (OD), object tracking (OT), semantic segmentation ($S$), motion prediction (MP) and accident prediction (AP).
@@
CODD [44] 2021
V2V
L OD 2 13k CARLA Scenario
V2X-Sim1.0 [45] 2021
V2V
L & RGBD OD 3 10k CARLA & Sumo Scenario
OPV2V [130] 2022
V2V
L & RGB OD 1 12k OpenCDA Scenario
V2X-Sim 2.0 [131] 2022
V2V & V2I
L & RGBD OD, OT, SS 3 10k CARLA & Sumo Scenario
V2XSet [16] 2022
V2V & V2I
L OD 1 12k OpenCDA Scenario
IRV2V [48] 2023
V2V
L & RGB OD NA 8k CARLA Scenario
OPV2V-H [77] 2024
V2V
L & RGBD OD 1 10k OpenCDA Scenario
Semantic OPV2V [96] 2024
V2V
L & RGBD $S$ 12 10k OpenCDA Scenario
DeepAccident [103] 2024
V2V & V2I
L & RGB OD, OT, SS, MP, AP 2 57k CARLA Weather, daytime
TABLE XXII: Overview of more technical aspects of the synthetic datasets. Default synchronization means perfect synchronization